please help me solve this problem

8-1 5/6
8 minus 1 and 5/6
it kind of slips my mind sometimes when it comes to subtracting fractions

Answer by Earlsdon(6294)

8 - 1 5/6
You can do this in two steps.
1) 8 - 1 = 7
2) 7 - 5/6 = 6 1/6
To make the second a little easier to understand, you borrow a 1 from the 7 leaving 6. Now you change the 1 you borrowed into 6/6, because 1 = 6/6. (7 = 6 + 6/6)
So now step 2) looks like this:
6&6/6 - 5/6 = 6&1/6
The answer:
8 - 1&5/6 = 6&1/6